Find the area, take the elements of the area parallel to the x-axis. y= 2x³-3x³-9x; y=x²-2x²-3x.

1
Expert's answer
2022-05-31T08:57:13-0400
"x^3-2x^2-3x=2x^3-3x^2-9x""x^3-x^2-6x=0""x(x^2-x-6)=0""x(x+2)(x-3)=0""x_1=-2, x_2=0, x_3=3""A_1=\displaystyle\int_{-2}^{0}(2x^3-3x^2-9x-(x^3-2x^2-3x))dx""=\displaystyle\int_{-2}^{0}(x^3-x^2-6x)dx""=[\dfrac{x^4}{4}-\dfrac{x^3}{3}-3x^2]\begin{matrix}
 0 \\
 -2
\end{matrix}"





"=0-(\dfrac{(-2)^4}{4}-\dfrac{(-2)^3}{3}-3(-2)^2)"





"=\dfrac{16}{3}({units}^2)""A_2=\displaystyle\int_{0}^{3}(x^3-2x^2-3x-(2x^3-3x^2-9x))dx""=\displaystyle\int_{0}^{3}(-x^3+x^2+6x)dx""=[-\dfrac{x^4}{4}+\dfrac{x^3}{3}+3x^2]\begin{matrix}
 3 \\
 0
\end{matrix}"





"=-\dfrac{(3)^4}{4}+\dfrac{(3)^3}{3}+3(3)^2-0"





"=\dfrac{63}{4}({units}^2)"